### v1

@@
Abstract
Scaling up language models to billions of parameters has opened up possibilities for in-context learning, allowing instruction tuning and few-shot learning on tasks that the model was not specifically trained for. This has achieved breakthrough performance on language tasks such as translation, summarization, and question-answering. Furthermore, in addition to these associative “System 1” tasks, recent advances in Chain-of-thought prompt learning have demonstrated strong “System 2” reasoning abilities, answering a question in the field of artificial general intelligence whether LLMs can reason.
@@ Abstract
The field started with the question whether LLMs can solve grade school math word problems. This paper reviews the rapidly expanding field of prompt-based reasoning with LLMs. Our taxonomy identifies different ways to generate, evaluate, and control multi-step reasoning. We provide an in-depth coverage of core approaches and open problems, and we propose a research agenda for the near future. Finally, we highlight the relation between reasoning and prompt-based learning, and we discuss the relation between reasoning, sequential decision processes, and reinforcement learning. We find that self-improvement, self-reflection, and some metacognitive abilities of the reasoning processes are possible through the judicious use of prompts. True self-improvement and self-reasoning, to go from reasoning with LLMs to reasoning by LLMs, remains future work.
1 Introduction
Transformer-based Large Language Models (LLMs) that are trained on large datasets have achieved breakthrough performance at next token prediction (Vaswani et al., 2017; Radford et al., 2019; Wei et al., 2022a); they are very good at natural language understanding (GLUE, SQUAD, Xsum) (Wang et al., 2018, 2019; Rajpurkar et al., 2016; Narayan et al., 2018), translation (Kocmi et al., 2022; Papineni et al., 2002; Sennrich et al., 2015), question answering (Tan et al., 2023), and other System 1 tasks (Kahneman, 2011).111In his book Thinking, fast and slow, a bestseller on human psychology, Daniel Kahneman described System 1 thinking as a near-instantaneous process; it happens automatically, intuitively, and with little effort. It is driven by instinct and experiences. System 2 thinking is slower and requires more effort. It is conscious and logical. The automatic operations of System 1 generate surprisingly complex patterns of ideas, but only the slower System 2 can construct thoughts in an orderly series of steps. In the LLM literature the terms are often used as shorthand to distinguish single-step associative tasks, from multi-step reasoning tasks, despite the fact that language tasks such as question answering and translation may require some “slow” thinking. The success of ChatGPT (Ouyang et al., 2022) has taken the world by storm.
@@ 1 Introduction
Transformer-based generative language models whose size is beyond hundreds of billions parameters are not only very good at language generation, they also enable new type of machine learning, called in-context learning (Brown et al., 2020). In-context learning, also known as prompt-based learning, occurs only in LLMs beyond a certain size (hundreds of billions of parameters) that are sufficiently rich (Wei et al., 2022a). In-context learning is inference time, prompt-based, few-shot learning, where model parameters are not trained or fine-tuned.
System 1 tasks, such as associative language tasks, are easily solved by LLMs with prompt-based learning, as the many school children around the world that use ChatGPT daily can attest. (Although the problems are too often not solved correctly, just fluently, when the model’s association powers lead to hallucination (Huang et al., 2023).) On the other hand, System 2 tasks, such as grade school math word problems, are more difficult for LLMs(Cobbe et al., 2021). To solve math word problems we need to break down the problem in multiple reasoning steps. Spurred-on by the impressive performance on System 1 tasks, much research has focused on understanding the reason for the poor performance of LLMs on System 2 tasks, and how it can be improved.
Among this research, the Chain-of-thought experiment (Wei et al., 2022b) stands out. This work, and subsequently Kojima et al. (2022), showed that adding a simple instruction to the prompts, Let’s think step by step, can provoke an LLM to perform the required intermediate reasoning steps, achieving a surprising jump in performance. The Chain-of-thought paper is a breakthrough in the field of reasoning with LLMs. Much exciting work has been published that builds on this work.
@@
Grade school math word problems started the research into LLM-reasoning, with the GSM8K benchmark (Cobbe et al., 2021). In our survey we discuss papers based on this benchmark, and directly-related follow up work on reasoning. We focus on prompt-based approaches. We survey the recent literature using a straightforward taxonomy.
Although the field has only recently started, the jump in performance on reasoning has excited artificial intelligence and society alike. We provide a research agenda with opportunities for future research. At the end of this survey, we also discuss connections to other fields, such as self-reflection, metacognition (or thinking about thinking, see for example Dunlosky and Metcalfe (2008)), and the motivation towards artificial general intelligence.
Our contributions are:
A survey of relevant approaches in prompt-based reasoning (grade school math word problems and closely related domains) in large language models, including a research agenda.
A taxonomy based on regular reasoning literature (step generation, step evaluation, and control of reasoning steps).
@@
This survey is organized as follows. Section 2 summarizes the most relevant developments in LLMs, including in-context learning. Of great importance are the benchmarks that are used in this field. We discuss these in Section 3, followed by our method for scoping and selecting of papers in Section 4. Next, in Section 5 we provide a taxonomy of the field, where we discuss the approaches in detail. Then, in Section 6 we discuss our findings in a broader perspective. We also discuss the relation between reasoning and work on self-reflection and metacognition. This section concludes with an agenda for future research. Finally, Section 7 concludes the survey.
2 Background: Reasoning with LLMs
@@
Before we dive into the works on reasoning, we review some background terminology on LLMs. Our overview is brief. Excellent surveys on LLMs are, for example, Minaee et al. (2024) and Zhao et al. (2023). We discuss the generic training pipeline for LLMs, we discuss how in-context learning works, and we discuss the reasoning pipeline. We start with the generic language model training pipeline.
2.1 Training Pipeline Language Model
@@
LLMs are typically constructed in a sequence of stages, from data preparation, through training, to inference. The training pipeline for most LLMs is quite elaborate. We will now list a pipeline of the most common stages, based on the survey by Minaee et al. (2024).
1.
Acquire a large, general, unlabeled, high-quality text corpus. Some considerations on the selection of the texts are discussed in Brown et al. (2020).
2.
Pretrain the transformer model (Vaswani et al., 2017) on this large corpus. This step yields a generalist model. The pretraining is done using a self-supervised approach on the unlabeled dataset (text corpus).
3.
Finetune the general model to a specific (narrow) task. This can be done using supervised-learning with a new labeled dataset consisting of prompts and answers (supervised finetuning, SFT) (Wei et al., 2022a; Minaee et al., 2024), specific for the task at hand. (A small number of papers in this survey work in the finetuning stage.)
4.
Instruction tuning is a form of finetuning on a labeled dataset of instruction prompts and corresponding outputs, to improve instruction following, and thus the usefulness of models.
5.
Align the finetuned model with user expectations (preference alignment). The goal of this stage is to improve the model to give more ethically and socially acceptable answers. The machine learning method that is used in this stage can be, for example, Reinforcement Learning with Human Feedback (Ouyang et al., 2022) or Direct Preference Optimization (Rafailov et al., 2024).
- 6.
7.
Inference & In-context learning can be used to train the model to provide the correct answers without changing parameters (Dong et al., 2022; Brown et al., 2020). By providing a prompt that contains a small number of examples together with a question, prompt learning is a form of few-shot learning. This is the stage in which most of the papers of this survey work, and that is familiar to all general users of ChatGPT.
Most of the reasoning methods that we discuss in this survey work in stage 7: in-context learning, using prompts for the LLM to perform a complex multi-step reasoning task. The following section provides a brief introduction to in-context learning.
2.2 In-Context Learning
@@
In LLMs beyond hundreds of billions of parameters a new kind of learning has emerged, that is called in-context learning or prompt-learning (Brown et al., 2020). It occurs at inference time, and is often able to give good results with few examples; it is a form of few-shot learning. The large size of the model, containing rich and general knowledge is enabling this new type of few-shot learning (see Dong et al. (2022) for a survey).
@@
A prompt, consisting of a piece of demonstration context, is concatenated with a query question, and is given to the language model for prediction (Liu et al., 2023). For example, when the task is emotion recognition in a social media post, “I missed the bus today,” can be followed by “I felt so [___]”. Alternatively, for translation, we could follow “I missed the bus today,” by “French: [___]” (Liu et al., 2023). The prompt contains background information that is recognized by the model, selecting the desired model context. In-context learning works when language models contain enough knowledge, allowing them to generalize on the examples provided in the prompt.
Prompts that contain a few examples are said to perform few-shot learning. Prompts that contain only instructions without examples are said to perform zero-shot learning.
@@
In-context learning takes place at inference time, after the computationally intensive stages where parameters have been pretrained and finetuned, when the model is queried by the user to provide answers. No parameters are changed anymore with in-context learning. This is quite different from the common approach in supervised deep learning—or self-supervised deep learning—where large datasets are used during training to update model parameters with backward propagation in lengthy and costly training epochs (Goodfellow et al., 2016). Common approaches to few-shot learning, such as metalearning, do include training and finetuning of parameters to achieve generalization, and are computationally expensive (see, for example, Finn et al. (2017) or Huisman et al. (2021); Hospedales et al. (2021) for a survey).
@@
Prompts provide a user-friendly interface to LLMs. The success of in-context learning tends to be quite sensitive to the way a prompt is formulated; a new field called prompt engineering has emerged to optimize the usefulness of in-context learning by learning how to make them do what we want (Radford et al., 2019; Wei et al., 2022a; Giray, 2023; Sahoo et al., 2024).
2.3 Reasoning Pipeline
Reasoning problems are also solved with a pipeline of stages. A typical approach to solving a complex problem is to subdivide it into smaller steps and solve those. This approach is related to divide and conquer (Bellman, 1966). New steps are (1) generated, (2) evaluated, and the number of steps that are generated and searched is (3) controlled in some way. The in-context reasoning approaches that we survey follow a general three-stage pipeline (Madaan et al., 2023):
1.
Generate: generation of steps by the model,
2.
Evaluate: evaluation of the predicted steps by an evaluator,
3.
Control: control of the number of steps that are generated and how deep ahead the reasoning process will look.
This three-stage pipeline will be the basis of our taxonomy. But first, we will look at benchmarks.
3 Benchmarks
@@
Progress in artificial intelligence is measured by benchmarks. Benchmarks define the goal that researchers aim to achieve in their experiments. In natural language processing, a wide array of benchmarks exists to measure progress, such as on question answering (for example, CommonsenseQA (Talmor et al., 2018)), word prediction (for example, LAMBADA (Paperno et al., 2016)), translation (for example, WMT’22 (Kocmi et al., 2022)), language understanding (for example, GLUE (Wang et al., 2018, 2019)), and text summarization (for example, Xsum (Narayan et al., 2018)). Transformer architectures were first popularized by encoder models such as BERT (Devlin et al., 2018), for named entity recognition and classification tasks. Subsequently, decoder models such as GPT 2-4 (Radford et al., 2019; Brown et al., 2020; Achiam et al., 2023) showed impressive progress on natural language benchmarks.
The field of LLMs is quite active. Many different benchmarks exist, and listing a comprehensive overview of all relevant benchmarks is beyond the scope of this survey. We will mention relevant benchmarks for testing the reasoning abilities of LLMs. Following Wei et al. (2022b), these are all math word problem benchmarks. The benchmark that is most frequently associated with reasoning by LLMs is a dataset of grade school math word problems GSM8K (Cobbe et al., 2021). GSM8K was created by humans, with an aim of high quality, high diversity, moderate difficulty, and solutions in natural language. Other benchmarks are the SVAMP varying structures benchmarks (Patel et al., 2021), the ASDiv dataset of diverse math problems (Miao et al., 2021), the AQuA dataset of algebraic word problems (Ling et al., 2017), and the MAWPS benchmark (Koncel-Kedziorski et al., 2016).
We will now briefly discuss these benchmarks; the baseline performance that we quote is from Wei et al. (2022b).
GSM8K
To test reasoning skills, the Grade School Math problem dataset (GSM8K) was developed for testing LLMs (Cobbe et al., 2021). It consists of 8500 human-written math problems. Language models struggled to achieve good performance on this dataset (pre Chain-of-thought). An example of a math word task is:
Problem: Beth bakes 4, two dozen batches of cookies in a week. If these cookies are shared amongst 16 people equally, how many cookies does each person consume?
Answer: .
The baseline performance of GPT-3 175B is 15.6% accuracy. In comparison, the performance of Chain-of-thought is 46.9% accuracy.
ASDiv
The Academia Sinica Diverse MWP Dataset (ASDiv) (Miao et al., 2021) is specifically designed for high diversity in problem types, formats and difficulty levels. It consists of 2305 problems. An example problem is:
Problem: A sandwich is priced at . A cup of pudding is priced at . Tim bought 2 sandwiches and 4 cups of pudding. How much money should Tim pay?
Answer: .
The baseline performance of GPT-3 175B is 70.3% accuracy. The performance of Chain-of-thought is 71.3% accuracy.
MAWPS
The Math Word Problem Repository (MAWPS) (Koncel-Kedziorski et al., 2016) allows for the construction of datasets with particular characteristics by selecting different categories of problems. The dataset consists of 3320 problems. An example is:
Problem: Rachel bought two coloring books. One had 23 pictures and the other had 32. After one week she had colored 44 of the pictures. How many pictures does she still have to color?
Answer: .
The baseline performance of GPT-3 175B is 72.7% accuracy. The performance of Chain-of-thought is 87.1% accuracy.
SVAMP
The Simple Variations on Arithmetic Math word Problems dataset (SVAMP) was designed by Patel et al. (2021). It consists of 1000 problems, curated from variations of ASDiv-a (Miao et al., 2021) and MAWPS (Koncel-Kedziorski et al., 2016). An example problem is:
Problem: Jack had 8 pens and Mary had 5 pens. Jack gave 3 pens to Mary. How many pens does Jack have now?
Answer: .
The baseline performance of GPT-3 175B is 65.7% accuracy. In comparison, the performance of Chain-of-thought is 68.9% accuracy.
AQuA
The Algebraic Question Answering dataset (Ling et al., 2017) is a large dataset of 100,949 questions, answers, and rationales. The dataset is based on a combination of a smaller seed dataset and crowdsourcing. An example question is:
Question: Two trains running in opposite directions cross a man standing on the platform in 27 seconds and 17 seconds respectively and they cross each other in 23 seconds. The ratio of their speeds is: Options: A) 3/7 B) 3/2 C) 3/88 D) 3/8 E) 2/2
Answer: B.
The baseline performance of GPT-3 175B is 24.8% accuracy. The performance of Chain-of-thought is 35.8% accuracy.
There is a wide variety of benchmarks, and there is a wide variety of performance in benchmarks. Some are easily solvable by current LLMs, and some are significantly harder. Benchmark design is an important part of the field of reasoning in LLMs. Currently the GSM8K benchmark is popular; baseline model performance is weak, and reasoning prompts can substantially improve performance. As performance on GSM8K improves, different (harder) benchmarks will become popular.
4 Selection of Papers
The papers in this survey were selected as follows. Baseline LLMs have difficulty solving math word problems, specifically on benchmarks listed in the previous section. We take the ability to solve those benchmarks as a proxy for reasoning ability. We initially performed a literature search for papers that use these benchmarks, and that contain the search terms reasoning and large language model in their title or abstract. We also searched for papers that referenced the Chain-of-thought paper. The resulting papers were curated based on recency, relevance, substance, and novelty.
We favor recent papers (two years prior to the writing of the survey), related to the Chain-of-thought approach of generating intermediate reasoning steps, that solve tasks such as math word problems, and that work by prompt-based in-context learning. We also include some papers that work by finetuning or supervised learning that relate to, or inspire, the Chain-of-thought approaches. Furthermore, we include approaches outside math word problems that showed interesting approaches to reasoning, such as applications in coding and autonomous agents, because of their approach to grounding.
5 Prompt Generation, Evaluation and Control

### v2

@@
Abstract.
Language models with billions of parameters exhibit in-context learning abilities, enabling few-shot learning on tasks that the model was not specifically trained for. Traditional models achieve breakthrough performance on language tasks, but do not perform well on basic reasoning benchmarks. However, a new in-context learning approach, Chain-of-thought, has demonstrated strong multi-step reasoning abilities on these benchmarks.
@@ Abstract.
The research on LLM reasoning abilities started with the question whether LLMs can solve grade school math word problems, and has expanded to other tasks in the past few years. This paper reviews the field of multi-step reasoning with LLMs. We propose a taxonomy that identifies different ways to generate, evaluate, and control multi-step reasoning. We provide an in-depth coverage of core approaches and open problems, and we propose a research agenda for the near future.
We find that multi-step reasoning approaches have progressed beyond math word problems, and can now successfully solve challenges in logic, combinatorial games, and robotics, sometimes by first generating code that is then executed by external tools. Many studies in multi-step methods are using reinforcement learning for finetuning, external optimization loops, in context reinforcement learning, and self-reflection.
1. Introduction
Transformer-based Large Language Models (LLMs) that are trained on large datasets have achieved breakthrough performance at text generation tasks that directly build on next token prediction (Vaswani et al., 2017; Radford et al., 2019; Wei et al., 2022a); they are good at natural language understanding (GLUE, SQUAD, Xsum) (Wang et al., 2018, 2019; Rajpurkar et al., 2016; Narayan et al., 2018), translation (Kocmi et al., 2022; Papineni et al., 2002; Sennrich et al., 2016), question answering (Tan et al., 2023), and other language generation tasks. The success of models such as ChatGPT (Ouyang et al., 2022) is impressive.
@@ 1. Introduction
Transformer-based generative language models whose size is beyond hundreds of billions parameters are not only good at language generation, they also enable a new type of machine learning, called in-context learning (Brown et al., 2020). In-context learning, also known as prompt-based learning, is an emergent ability that occurs in LLMs beyond a certain size (hundreds of billions of parameters—less, with judicious prompting) that have been finetuned for conversational responses (Wei et al., 2022a). In-context learning is inference-time, prompt-based, few-shot learning with instructions. As opposed to finetuning, model parameters are not adapted by in-context learning.
Language generation tasks are solved well by LLMs with prompt-based learning. On the other hand tasks that require reasoning, such as grade school math word problems, are more difficult for LLMs (Cobbe et al., 2021). Spurred-on by the impressive performance on language tasks, much research has focused on understanding the reason for the poor performance of LLMs on reasoning tasks, and how it can be improved. Among this research, the Chain-of-thought paper by Wei et al. (2022b) stands out. This work, and later work by Kojima et al. (2022), showed that adding a simple instruction to the prompt, Let’s think step by step, can provoke an LLM to perform the required intermediate reasoning steps to answer difficult problems in a multi-step approach. Subsequently, performance on math word benchmarks has increased markedly. Much of the performance increase of models such as OpenAI o1, o3 (Huang et al., 2024) and DeepSeek R1 (Shao et al., 2024; Guo et al., 2025) is attributed to multi-step reasoning methods as reviewed here, with reinforcement learning playing an increasingly important role, both for in-context learning and finetuning (Xu et al., 2025; Wu et al., 2024a; Du et al., 2025; Chu et al., 2025).
@@
The line of research into multi-step LLM-reasoning was started with grade school math word problems, with the GSM8K benchmark (Cobbe et al., 2021). Soon, other reasoning domains were included, such as reasoning about advanced math problems, computer code, robotic movement, and games. Many works have been published that build on Chain-of-thought (Chu et al., 2024b). In this paper, we survey the literature using a straightforward taxonomy. We discuss papers based on several reasoning benchmarks, as well as directly-related follow up work.
Having started with basic math word problems, multi-step LLM reasoning approaches now perform logic reasoning, planning, combinatorial games, and robotic actions, amongst others. Some approaches do so by prompting the LLM to generate code that is then interpreted by external tools. Many studies in multi-step methods are using reinforcement learning, for finetuning, in context reinforcement learning, external optimization loops, and self-reflection. The main contributions of this paper are: (1) we provide a survey of relevant approaches in multi-step reasoning with LLMs, (2) we propose a taxonomy based on the reasoning literature (step generation, step evaluation, and control of reasoning steps), and (3) we formulate a research agenda for reasoning with LLMs.
@@
This survey is organized as follows. Section 2.1 provides background information on the most relevant developments in LLMs, including in-context learning and finetuning. Of great importance are the benchmarks that are used in this field (Section 2.4). Next, in Section 3 we provide a taxonomy of the field, where we discuss the main approaches in detail. Then, in Section 4 we discuss our findings in a broader perspective. We also discuss the relation between multi-step reasoning and work on self-reflection and metacognition. This section concludes with an agenda for future research. Finally, Section 5 concludes the survey.
2. Background, Scope and Selection of Papers
Reasoning has a long and active history in AI, in logical inference, and in other fields, such as commonsense and analogical reasoning (Lewis and Mitchell, 2024; Mitchell and Hofstadter, 1990). Indeed, some of the early criticism on LLMs was that they could not reason, and that they made obvious and basic reasoning errors, showing a lack of understanding. Berglund et al. (2024) provide the example of a model that is trained to report that “Valentina Tereshkova was the first woman to travel to space,” but is not able to answer the question, “Who was the first woman to travel to space?” pointing indeed to a lack of semantic understanding. Other work suggests that results are less generalizable and transferable than often assumed, showing how base-10 arithmetic skills do not transfer to base-9 arithmetic problems (Wu et al., 2024b). Different kinds of reasoning have played an important role in LLM research.
2.1. Scope
The interest in Chain-of-though has stimulated more research into language and logic. Logic-LM combines LLMs with symbolic logic solvers in an agentic way, where the symbolic solver is called by the LLM as a tool (Pan et al., 2023). Xu et al. (2024a) use a similar approach. Relevant logic-reasoning benchmarks are PrOntoQA (Saparov and He, 2023), ProofWriter (Tafjord et al., 2020), FOLIO (Han et al., 2022), LogicalDeduction (Srivastava et al., 2022), and AR-LSAT (Zhong et al., 2022). The logic-LLM approaches often combine Chain-of-thought and symbolic solvers in an agentic-LLM approach (Plaat et al., 2025), or transform the problem in an intermediate formal language. The use of LLMs for generating formal languages is further discussed in section 4.4.
More recently, LLMs are also shown to struggle with analogies and analogical reasoning (Lewis and Mitchell, 2024; Mitchell and Hofstadter, 1990), a topic that is related to self-reflection (see Sections 3.3.3 and 4.3.3).
The question whether LLMs can reason prompted the development of the GSM8k benchmark (Cobbe et al., 2021) in 2021, a benchmark of some of easy reasoning problems (of the type: “Question: Natalia sold clips to 48 of her friends in April, and then she sold half as many clips in May. How many clips did Natalia sell altogether in April and May? Answer: 72”). This benchmark was specifically designed to test if LLMs can answer basic mathematical multi-step reasoning questions, or if they can not reason at all. Initially, performance was poor, but a year later Wei et al. (2022b) showed that by providing multi-step examples in the prompt—a chain of thought—much better performance was possible. Many works followed, and reasoning in language models has become an active area of study. In this survey, we focus on prompt-based multi-step reasoning algorithms. After starting with math word problems, work has more recently been extended to combinatorial games, puzzles, robotics, logic, and other fields of reasoning. At the end of this survey, we discuss connections to other fields, such as self-reflection and in-context reinforcement learning.
Before we discuss the works on reasoning, this section reviews background terminology on LLMs. Our overview is brief. Excellent recent general surveys on LLMs are, for example, Minaee et al. (2024) and Zhao et al. (2023). Other works focus on the nature of reasoning and its definition (Yu et al., 2023; Chu et al., 2024a; Huang and Chang, 2023), on evaluating logical reasoning in LLMs (Mondorf and Plank, 2024; Liu et al., 2023a; Pan et al., 2023), on reinforcement learning in reasoning (Xu et al., 2025), on reasoning in language (Yu et al., 2023), or on Chain-of-thought itself (Chu et al., 2024b).
The papers in this survey were selected as follows. We started by selecting papers on the ability to solve math word benchmarks (as a proxy for reasoning ability), that contained the search terms reasoning and large language model in their title or abstract, with a focus on papers that reference the Chain-of-thought paper. Although multi-step reasoning with LLMs initially was aimed at solving math world problems, it is now wider, including benchmarks and approaches for computer code, game play, puzzles, robot movement, and webpage navigation (see Table 2). We selected recent papers (two years prior to the writing of the survey) that show experimental results on selected benchmark datasets.
We focus on prompt-based, in-context learning, methods based on Chain-of-thought, that are used in reasoning LLMs such as OpenAI o1 and o3 (Xu et al., 2025; Wu et al., 2024a; Huang et al., 2024). We also include papers that work with external algorithms, finetuning or supervised learning that have contributed to the approaches.
@@ 2.1. Scope
We discuss the generic training pipeline for LLMs, we discuss how in-context learning works, and we discuss commonly used benchmarks. We start with the generic language model training pipeline.
2.2. Language Model Training Pipeline
@@
LLMs are typically constructed in a sequence of stages, from data preparation, through training, to inference. The training pipeline for most LLMs is quite elaborate. We will now list a brief pipeline of the most commonly used stages, based on the survey by Minaee et al. (2024).
In training an LLM, the first stage is to acquire a large, general, unlabeled, high-quality text corpus. Some considerations on the selection of the texts are discussed in Brown et al. (2020). The next stage is pretraining the transformer model (Vaswani et al., 2017) on this large corpus. This stage yields a generative language model. Pretraining is done using self-supervised autoregressive training on the unlabeled dataset (text corpus). Then the general model is finetuned to a specific (narrow) task. This can be done using supervised learning with a new labeled dataset consisting of prompts and answers (supervised finetuning, SFT, for example with low-rank optimization (Hu et al., 2022; Wei et al., 2022a; Minaee et al., 2024), or reinforcement learning (Sutton and Barto, 2018; Chu et al., 2025), specific for the task at hand. A small number of the surveyed papers also address the finetuning stage. A specific form of finetuning is instruction tuning, to improve instruction following for a certain task. Instruction tuning is supervised by a labeled dataset of instruction prompts and corresponding outputs. Depending on the purpose of the model the next step is alignment of the finetuned model with user expectations (preference alignment). Alignment is usually performed to ensure that the model produces more ethically and socially acceptable answers, preventing, for example, hate speech. The machine learning method that is commonly used in this stage Reinforcement Learning with Human Feedback (Ouyang et al., 2022) or Direct Preference Optimization (Rafailov et al., 2024). Optionally, model training can be optimized to improve cost-effectiveness, for example, mixed precision training (Micikevicius et al., 2018), quantization (Jacob et al., 2018), or knowledge distillation (Xu et al., 2024b; Gu et al., 2023).
Once the model has been trained in the steps described above, it can be used further in the inference stage. Here, the model is used to provide an answer to the prompt. The inference-time stage is post-training, no model parameters are changed anymore (Dong et al., 2023; Brown et al., 2020); in-context learning, or prompt-learning, takes place in this stage. Due to its low barrier of entry, without the need for expensive training or finetuning, this form of advanced prompt engineering has become a popular training method. This is the stage on which most of the surveyed papers focus, using prompts for the LLM to perform a complex multi-step reasoning task. The following section provides a brief introduction to in-context learning.
2.3. In-Context Learning
@@
In large models, beyond hundreds of billions of parameters, a new kind of learning has emerged, that has been called in-context learning or prompt-learning (Brown et al., 2020). It occurs not when the model is trained, but when it is used, at inference time. Since no parameters are changed in this stage, it is not a model training stage; in-context-learning “learns” inside the context, or prompt, using information that is already encoded in the trained model parameters and the prompt, not by training the model anymore. In-context learning is often able to give good results with few examples, so-called few-shot learning, learning from the few examples in combination with the knowledge of the model. The large size of the model, containing rich and general knowledge, is enabling the few-shot learning (see Dong et al. (2023) for a survey).
@@
In in-context learning, a prompt, consisting of a piece of demonstration context, is concatenated with a query question, and is given to the language model, for text generation (Liu et al., 2023b). For example, when the task is emotion recognition in a social media post, “I missed the bus today,” can be followed by “I felt so [___]”, and the model could answer with “bad”. Alternatively, for translation, we could follow “I missed the bus today,” by “French: [___]” to request a translation (Liu et al., 2023b). The prompt contains background information that is recognized by the model, selecting the desired model context. In-context learning works when language models contain enough knowledge, allowing them to generalize on the (few) examples provided in the prompt.
@@
Contexts that contain a few examples are said to perform few-shot learning. Contexts that contain only instructions with zero examples are said to perform zero-shot learning. In-context learning takes place at inference time, after the computationally intensive training stages where parameters have been pretrained and finetuned, when the model is queried by the user to provide answers. No parameters are changed anymore with in-context learning. This is quite different from the common approach in supervised deep learning—or self-supervised deep learning—where large datasets are used during training to update model parameters with backward propagation in lengthy and costly training epochs (Goodfellow et al., 2016). Indeed, in-context learning takes place fully at inference time, no parameters are trained, instead, learning now refers to adjusting the answers to the examples in the prompt and the internal knowledge acquired during training. Common approaches to few-shot learning, such as metalearning, do include training and finetuning of parameters to achieve generalization, and are computationally expensive (see, for example, (Finn et al., 2017) or (Huisman et al., 2021; Hospedales et al., 2021) for a survey). In-context learning, in comparison, is computationally cheap, and it has become a popular research approach..
@@
Prompts provide a user-friendly interface to LLMs. The success of in-context learning tends to be quite sensitive to the way a prompt is formulated; a new field called prompt engineering has emerged to help human users learn how to make LLMs do what we want them to do (Radford et al., 2019; Wei et al., 2022a; Giray, 2023; Sahoo et al., 2024). The current survey thus discusses advanced prompt engineering methods.
2.4. Multi-step Reasoning Benchmarks
@@
Progress in artificial intelligence is measured by benchmarks. Benchmarks define the goal that researchers aim to achieve in their experiments. In natural language processing, a wide array of benchmarks exist to measure progress, such as on question answering (for example, CommonsenseQA (Talmor et al., 2019)), word prediction (for example, LAMBADA (Paperno et al., 2016)), translation (for example, WMT’22 (Kocmi et al., 2022)), language understanding (for example, GLUE (Wang et al., 2018, 2019)), and text summarization (for example, Xsum (Narayan et al., 2018)).
The field of LLMs is quite active. We will mention relevant benchmarks for testing the multi-step reasoning abilities of LLMs. The research on reasoning with LLMs started with math word problem benchmarks. The benchmark that is most frequently associated with multi-step reasoning with LLMs is the dataset of grade school math word problems GSM8K (Cobbe et al., 2021). GSM8K was created with the aim of providing high quality, high diversity, moderate difficulty, problems and solutions in natural language. It consists of 8500 human-written math problems. Language models struggled to achieve good performance on this dataset before Chain-of-thought was introduced. An example of a math word task follows. Problem: Beth bakes 4 trays with two dozen batches of cookies in a week. If these cookies are shared amongst 16 people equally, how many cookies does each person consume? Answer: .
Other benchmarks of similar math word problems are the SVAMP varying structures benchmarks (Patel et al., 2021), the ASDiv dataset of diverse math problems (Miao et al., 2020), the AQuA dataset of algebraic word problems (Ling et al., 2017), and the MAWPS benchmark (Koncel-Kedziorski et al., 2016). Table 1 summarizes the accuracy of Chain-of-thought on these basic math word problems, against the baseline of GPT-3 175B as LLM (Wei et al., 2022b), as percentage of benchmark questions answered correctly.
| Benchmark | GPT-3 175B | Chain-of-thought (Wei et al., 2022b) |
| GSM8K | 15.6 | 46.9 |
| ASDiv | 70.3 | 71.3 |
| MAWPS | 72.7 | 87.1 |
| SVAMP | 65.7 | 68.9 |
| AQuA | 24.8 | 35.8 |
We see that Chain-of-thought performs well against the baseline of GPT-3 on some benchmarks, but there is certainly room for further improvement on others.
In addition to the initial set of math word benchmarks, further reasoning approaches have been introduced that test performance in other fields of reasoning. Benchmarks have been developed on Advanced mathematical questions (Codeforces, ), Computer code comprehension (Human evaluation, Spider (Yu et al., 2018), Transcoder (Roziere et al., 2020)), Robotic movement (Alfworld (Shridhar et al., 2021), Kitchen (Ahn et al., 2022)), Puzzle solving (Game24 (Yao et al., 2024)), Creative writing (Yao et al., 2024)), Gaming (Checkmate problems(Yang et al., 2024), MineCraft (Fan et al., 2022)), and Webpage navigation (WebShop (Yao et al., 2022)). These benchmarks are used by other approaches in our survey, as we will see in more detail in Section 3.
The scope of this survey is limited to multi-step reasoning approaches. Other studies have published more challenging benchmarks, to study the performance of Chain-of-thought and self-reflection in LLMs in puzzles and (logic) games (Ruoss et al., 2024; Paglieri et al., 2024; Wen et al., 2025).
It can be difficult to directly compare different benchmarking results, since there may be differences in the way the experiments were conducted. Kamoi et al. (2024) provide a critical analysis of benchmarking problems in LLM reasoning. Despite these challenges, general conclusions can be drawn. Section 4.1 provides guidance on how to choose reasoning approaches for different problem types.
3. Step Generation, Evaluation and Control